# Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer

### v1

@@ 4.1 Differentiable Non-uniform INT2 Group Quantizer
where $u(x)$ is the step function and transition point $t_{i}$ (i.e., the right edge of each quantization bin) is set to the center of each partition, computed as $t_{1}=p_{1}/2$ and $t_{i}=t_{i-1}+(p_{i-1}+p_{i})/2$. The straight-through estimator is applied to all step functions so that every parameter (including LLM weights, clipping, and partitioning parameters) can be updated via backpropagation.
@@
4.2 Dequantization for Non-Linear Weight Partitions
Unlike the usual uniform quantization scheme, mapping the quantized weights back to real values is not trivial in NU quantizers as the design space of NU dequantization method is large and the inference pipeline is directly affected. In NU2U Liu et al. (2022), the quantized weight $\textbf{W}_{q}$ is simply dequantized to a uniform grid. Similarly, LLT Wang et al. (2022) trains a learnable lookup table for weight and activation to adjust the quantization bin widths but keeps the dequantization the same as other common uniform quantizers.
This design choice has an obvious advantage: Inference can be done on existing hardware without any modification. However, we propose non-uniform dequantization as described in Eqn. 6, based on our observation that uniform dequantization can lead to performance drop under extremely low bit configurations, especially on smaller models.
$\begin{split}\textbf{W}_{deq}=&\sigma(\beta)\text{min}(\textbf{W})\\ +\frac{h}{3}\biggl{(}&\text{u}(\frac{\textbf{W}}{h}-t_{1})(w_{1}-w_{0})\\ +&\text{u}(\frac{\textbf{W}}{h}-t_{2})(w_{2}-w_{1})\\ +&\text{u}(\frac{\textbf{W}}{h}-t_{3})(w_{3}-w_{2})\biggr{)}\end{split}$ (6)
The procedure is designed upon Eqn. 5 with several modifications. The value range is shifted and scaled from $[0,1]$ to $[\sigma(\beta)\text{min}(\textbf{W}),\sigma(\gamma)\text{max}(\textbf{W})]$. The dequantization grid $w_{i}$ is defined in Eqn. 7; the first and last values are the minimum and maximum values of the normalized weight and the middle values are set to the center of the two consecutive transition points.
$w_{i}=\begin{cases}0,&\text{if }i=0\\ \frac{t_{i}+t_{i+1}}{2},&\text{if }0<i<3\\ 1,&\text{if }i=3\end{cases}$ (7)
The dequantization LUT for a quantization unit (a weight group of size 128 in this paper’s experiments) can be pre-computed without any runtime overhead as follows:
$\hat{\textbf{W}}=\{\hat{\text{W}}_{0},\hat{\text{W}}_{1},\hat{\text{W}}_{2},\hat{\text{W}}_{3}\}$ (8)
where $\hat{\text{W}}_{i}=\sigma(\beta)\text{min}(\textbf{W})+h\cdot w_{j}$.
4.3 An NF3 Variant of LDP
We apply not only 2-bit weight quantization but also 3-bit quantization using the asymmetric NF format of AFPQ Zhang et al. (2023) where separate scale values are computed for the negative and positive weights ($s_{neg}=\text{max}(|\textbf{W}_{neg}|)$, $s_{pos}=\text{max}(\textbf{W}_{pos})$). Although shown effective, such NF3 quantizer can lead to suboptimal performance when the distribution is not zero-centered. Therefore, we make a further improvement by applying the proposed LDP to this situation.
The idea is to employ the same learnable clipping parameters ($\beta$, $\gamma$) to obtain the quantization range $h$ and one partitioning parameter $s_{1}$ to express the learnable center point as $c=\sigma(\beta)\text{min}(\textbf{W})+h\cdot\sigma(s_{1})$. Then, the two scale values are updated as follows:
$\begin{split}s_{neg}=|c-\sigma(\beta)\text{min}(\textbf{W})|,\\ s_{pos}=|\sigma(\gamma)\text{max}(\textbf{W})-c|,\end{split}$ (9)
and the quantization process is derived as follows:
$\textbf{W}_{q}=\begin{cases}\lfloor\frac{\textbf{W}-c}{s_{pos}}\rceil,&\text{if }\textbf{W}>c\\ \lfloor\frac{\textbf{W}-c}{s_{neg}}\rceil,&\text{otherwise.}\end{cases}$ (10)
4.4 W2A4 Look-up Table (LUT) Inference
Designing the inference pipeline of non-uniform W2A4-quantized models poses a big challenge. First, efficient INT tensor cores cannot be utilized since accumulating the multiplication results in INT quantized space makes it impossible to dequantize the weights back to correct non-uniform real values in the LUT $\hat{\textbf{W}}$. Second, both weights and activations must undergo online dequantization to support dynamic quantization, which adds a large amount of computation overhead.
Therefore, we focus on designing GEMV kernel for LUT decoding predominantly bounded by memory bandwidth, which is ideal for featuring the advantage of extreme W2A4KV4 compression. We report our early exploration on GEMM design in Section A.5.
Kernel Arguments and Block Tiling
@@ 4.4 W2A4 Look-up Table (LUT) Inference
Each thread block consists of 128 threads (4 warps) and we only tile along the output dimension and define the tile size as BH. The reason we do not follow the traditional 2-dimensional tiling is that both the input tokens and weights are stored in row-major format and have sub-byte packing along the column direction, which makes it hard to efficiently use high-bandwidth memory that performs best when reading 128B data consecutively. Also, global loads with small transactions and repeated shared stores complicate the pipeline design for latency hiding and degrade overall performance.
Data Path and Latency Hiding Pipeline
@@
#Bits (W-A-KV) Configuration LLaMA-1 7B LLaMA-2 7B LLaMA-3 8B LLaMA-3.2 1B LLaMA-3.2 3B
Method Rotation LDP MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki
16-16-16 35.10 68.40 5.68 46.45 61.67 5.47 68.40 72.93 6.10 32.20 58.90 9.74 58.00 65.30 7.82
2-4-16 BitDistiller 25.88 42.56 23.19 26.24 43.36 16.47 23.11 39.46 Inf 25.00 36.82 Inf 24.41 37.89 Inf
Ours 26.75 52.28 8.79 26.04 51.49 8.93 29.80 50.59 13.68 25.00 41.08 31.32 29.60 45.29 18.79
Ours 26.98 52.46 8.28 28.04 51.10 8.18 31.87 50.86 12.48 26.30 41.35 27.46 31.40 45.71 16.96
2-4-4 BitDistiller 24.45 43.08 19.98 26.59 44.93 17.40 23.29 39.75 Inf 24.66 37.55 Inf 24.62 37.26 Inf
Ours 26.98 52.21 8.92 26.41 51.10 8.93 29.66 49.80 14.05 24.74 40.77 33.86 31.44 44.26 19.58
Ours 27.34 52.29 8.28 26.92 51.22 8.31 31.01 50.41 12.69 25.62 41.80 29.30 30.33 45.56 17.52
3-4-16 BitDistiller 26.88 55.68 7.47 31.72 56.15 7.04 42.24 55.39 10.19 26.06 37.53 Inf 25.22 37.32 Inf
Ours 28.70 58.52 6.44 34.30 59.28 6.25 54.16 61.06 7.92 26.45 47.88 13.75 47.34 55.66 9.82
Ours 29.46 59.39 6.39 37.33 59.74 6.23 55.33 61.53 7.80 27.77 48.18 13.68 47.31 55.87 9.74
3-4-4 BitDistiller 27.04 56.05 7.54 30.19 55.51 7.15 40.70 56.35 10.46 25.48 38.75 Inf 25.91 37.27 Inf
Ours 28.80 58.48 6.45 33.46 58.53 6.36 51.74 59.69 8.04 26.11 47.14 14.58 46.08 55.08 10.05
Ours 30.00 58.55 6.39 36.07 59.27 6.33 52.55 61.11 7.95 26.54 47.71 14.44 46.40 55.12 9.99
Table 1: Comparison of the perplexity score on WikiText2, MMLU (5s), and 0-shot. 0-shot is average score of 4 Zero-shot Common Sense Reasoning tasks. We show the perplexity results >100 by Inf. Full results of Zero-shot tasks are in the Appendix.
@@ 5.1 Experimental Settings
We evaluate RCP on LLaMA-1 Touvron et al. (2023a) 7B, LLaMA-2 Touvron et al. (2023b) 7B, LLaMA-3 AI@Meta (2024)(1B, 3B, 8B). Our evaluation of RCP was carried out on PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-c Clark et al. (2018) and MMLU Hendrycks et al. (2020). We use LLM-Humaneval-Benchmarks Chen et al. (2021) and GSM8K Cobbe et al. (2021) for reasoning tasks evaluation. We also report the perplexity score on WikiText2 Merity et al. (2016) for our evaluation.
@@
Experiment Configuration
[…]
@@ 5.2 Results
The results are summarized in Table 1. From the perspective of general language tasks, our method demonstrates the ability to quantize activations and KV-cache under the W2 settings to 4-bit, which was previously unattainable using existing QAT methods. The application of rotation effectively addresses the outlier issues, a common bottleneck in quantization, enabling stable performance even in extremely low-bit quantization scenarios. Furthermore, the addition of LDP not only improves performance on general language tasks but also enhances the accuracy of zero/few shot tasks, which were not adequately addressed by rotation alone. In case of LLaMA-2 7B W2A4KV4, a performance degradation is observed when using rotation only compared to the baseline. However, by incorporating LDP, consistent performance improvements were achieved.
@@
For the coding domain-specific model, WizardCoder Luo et al. (2023), BitDistiller failed to offer the functional quantized models in both W3 and W2 settings. In our method, applying rotation alone was not effective in W2 settings and recovered some output quality in W3 settings. By incorporating LDP, we achieved up to a threefold improvement in performance, with accuracy increasing from 6.09% to 23.20% under the W2A4KV4 configuration. As shown in Fig. 8 with the application of LDP, we were able to produce logically correct code outputs and eliminate repetition of meaningless code generation.
For the mathematical reasoning model, MetaMath Yu et al. (2023), the baseline BitDistiller and ours without LDP failed to offer functional quantized models while ours with LDP could produce working quantized models. These results highlight the critical role of LDP in enabling proper task performance for reasoning models under extreme low-bit quantization. The output comparison for this task is summarized in Fig. 2.
@@
Table 3 and 4 present the results for GEMV in terms of latency and memory comsumption. The latency of GEMV, excluding the activation quantization overhead, is faster compared to FP16 and QuaRot Ashkboos et al. (2024b). This improvement can be attributed to the lower bit precision, which enhances computational efficiency. Table 4 measures the peak memory footprint for W2A4 and W3A4. For W2A4, a significant reduction on 5.29x in memory footprint was achieved compared to FP16. Note that in the LLaMA-3.2 series, it is necessary to separate the embedding table and head modules to satisfy the invariance arising from their tying. Furthermore, as the size of the embedding table has increased compared to previous models, the compression ratio has decreased accordingly.

### v2

@@ 4.3 Partition: Learnable Direct Partitioning
Finally, when computing the loss function in Eqn. 3, each weight $\mathbf{\theta}_{S}\in\Theta_{S}$ is fake-quantized by Eqn. 10 as $\mathbf{\theta}_{S}\leftarrow DQ_{LDP}(\mathbf{\theta}_{s})$. We note that during the fake quantization, every step function $u(\cdot)$ is applied with the straight-through estimator so that every parameter (including LLM weights, clipping, and partitioning parameters) can be updated via backpropagation.
Application of LDP on NF3 format
We apply not only 2-bit integer weight quantization but also 3-bit quantization using the asymmetric NF format of AFPQ Zhang et al. (2023) where separate scale values are computed for the negative and positive weights ($s_{neg}=\text{max}(|\textbf{W}_{rc,neg}|)$, $s_{pos}=\text{max}(\textbf{W}_{rc,pos})$). Although shown to be effective, such an NF3 quantizer can lead to suboptimal performance when the distribution is not zero-centered. Therefore, we make a further improvement by applying the LDP to this situation.
The idea is to employ the same learnable clipping parameters ($\beta$, $\gamma$) to obtain the dynamic quantization range $h_{lwc}$ and one partitioning parameter $s_{1}$ to express the learnable center point as $c=\sigma(\beta)\text{min}(\textbf{W})+h\cdot\sigma(s_{1})$. Then, the two scale values are updated as follows:
$\begin{split}s_{neg}=|c-\sigma(\beta)\text{min}(\textbf{W}_{rc})|,\\ s_{pos}=|\sigma(\gamma)\text{max}(\textbf{W}_{rc})-c|,\end{split}$ (12)
and the quantization process is derived as follows:
$\textbf{W}_{q}=\begin{cases}\lfloor\frac{\textbf{W}_{rc}-c}{s_{pos}}\rceil,&\text{if }\textbf{W}_{rc}>c\\ \lfloor\frac{\textbf{W}_{rc}-c}{s_{neg}}\rceil,&\text{otherwise.}\end{cases}$ (13)
@@
4.4 Look-up Table (LUT) Inference for Non-uniform W2A4 GEMV
To implement the non-uniform W2A4 inference on modern GPUs, LUT-based GEMM and GEMV kernels are designed. The quantized INT2 weights $\mathbf{W_{q}}$ and the FP16 dequantization LUT $\hat{\mathbf{W}}$ are pre-computed from LDP using Eqn. 9 and 10 without incurring any runtime overhead as follows:
$\begin{split}\mathbf{W}_{q}=Q_{LDP}(\mathbf{W}_{rc}),\\ \hat{\textbf{W}}=\{\hat{\text{W}}_{0},\hat{\text{W}}_{1},\hat{\text{W}}_{2},\hat{\text{W}}_{3}\}\\ \mathrm{where}\;\hat{\text{W}}_{i}=\sigma(\beta)\mathrm{min}(\mathbf{W}_{rc})+h_{lwc}\cdot w_{i}.\end{split}$ (14)
However, designing such kernels poses a big challenge. First, efficient INT tensor cores cannot be utilized since accumulating the multiplication results in INT quantized space makes it impossible to dequantize the weights back to correct non-uniform real values in the LUT $\hat{\textbf{W}}$. Second, both weights and activations must undergo online dequantization to support dynamic quantization, which adds a large amount of computation overhead.
Therefore, we focus on designing GEMV kernel for LUT decoding predominantly bounded by memory bandwidth, which is ideal for featuring the advantage of extreme W2A4KV4 compression. We report our exploratory results on GEMM design in Section A.6.
Fast GEMV via Latency Hiding
@@ 4.4 Look-up Table (LUT) Inference for Non-uniform W2A4 GEMV
#Bits (W-A-KV) Configuration LLAMA-1 7B LLAMA-2 7B LLAMA-2 13B LLAMA-3 8B LLAMA-3.2 1B LLAMA-3.2 3B
Method Rotation LDP MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki
16-16-16 BF16 35.10 68.40 5.68 46.45 61.67 5.47 55.54 63.02 4.88 68.40 72.93 6.10 32.20 58.90 13.40 58.00 65.30 10.70
2-4-16 BitDistiller 25.88 42.56 23.19 26.24 43.36 16.47 26.05 39.66 23.16 23.11 39.46 Inf 25.00 36.82 Inf 24.41 37.89 Inf
BitDistiller 26.75 52.28 8.79 26.04 51.49 8.93 29.97 48.48 7.55 29.80 50.59 13.68 25.00 41.08 31.32 29.60 45.29 18.79
RCP 27.34 52.29 8.28 28.04 51.10 8.18 37.27 51.71 7.27 31.87 50.86 12.48 26.30 41.35 27.46 31.40 45.71 16.96
2-4-4 BitDistiller 24.45 43.08 19.98 26.59 44.93 17.40 24.72 36.73 32.43 23.29 39.75 Inf 24.66 37.55 Inf 24.26 37.26 Inf
BitDistiller 26.98 52.21 8.92 26.41 51.10 8.93 24.18 43.55 11.45 29.66 49.80 14.05 24.74 40.77 33.86 31.44 44.26 19.58
RCP 27.34 52.29 8.28 26.92 51.22 8.31 35.49 48.18 7.95 31.01 50.41 12.69 25.62 41.80 29.30 30.33 45.56 17.52
3-4-16 BitDistiller 26.88 55.68 7.47 31.72 56.15 7.04 42.68 54.59 6.99 42.24 55.39 10.19 26.06 37.53 Inf 25.22 37.32 Inf
BitDistiller 28.70 58.52 6.44 34.30 59.28 6.25 46.91 58.99 5.62 54.16 61.06 7.92 26.45 47.88 13.75 47.34 55.66 9.82
RCP 29.46 59.39 6.39 37.33 59.74 6.23 50.84 60.52 5.49 55.33 61.53 7.80 27.77 48.18 13.68 47.31 55.87 9.74
3-4-4 BitDistiller 27.04 56.05 7.54 30.19 55.51 7.15 40.58 54.57 9.02 40.70 56.35 10.46 25.48 38.75 Inf 25.91 37.27 Inf
BitDistiller 28.80 58.48 6.45 33.46 58.53 6.36 47.86 58.78 6.06 51.74 59.69 8.04 26.11 47.14 14.58 46.08 55.08 10.05
RCP 30.00 58.55 6.39 36.07 59.27 6.33 48.47 58.83 5.57 52.55 61.11 7.95 26.54 47.71 14.44 46.40 55.12 9.99
Table 2: Comparison of the perplexity score on WikiText2, MMLU (5s) and averaged accuracy on four Zero-shot Common Sense Reasoning tasks. We show the perplexity results >100 by Inf. Full results of Zero-shot tasks are in the Appendix.
@@ 5.1 Experimental Settings
We evaluate RCP on LLaMA-1 Touvron et al. (2023a) 7B, LLaMA-2 Touvron et al. (2023b)(7B, 13B), LLaMA-3 AI@Meta (2024)(1B, 3B, 8B). Our evaluation of RCP was carried out on PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-c Clark et al. (2018), MMLU Hendrycks et al. (2020) and LongBench Bai et al. (2024). We use LLM-Humaneval-Benchmarks Chen et al. (2021) and GSM8K Cobbe et al. (2021) for reasoning task evaluation. We also report the perplexity score on WikiText2 Merity et al. (2016) for our evaluation.
@@
Training Configurations
We set the weight learning rate to 8e-7 for WllA4 and 1e-6 for WllA4KV4, while the learning rate for LWC and LDP was set to 1e-5. We set the training sequence length to 1024 and the evaluation sequence length to 2048.
@@ 5.2 Results
We compare our proposed RCP with the state-of-the-art QAT method, BitDistiller Du et al. (2024). Details on training cost and implementation are provided in Appendix A.4.
@@
The results are summarized in Table 2. From the perspective of general language tasks, our method demonstrates the ability to quantize activations and KV-cache under the W2 settings to 4-bit, which was previously unattainable using existing QAT methods. The application of rotation effectively addresses the outlier issues, a common bottleneck in quantization, enabling stable performance even in extremely low-bit quantization scenarios. Furthermore, the addition of LDP improves performance on general language tasks across the board, and generally enhances the accuracy of zero/few-shot tasks, which were not adequately addressed by rotation alone. For example, the addition of LDP contributes to a performance gain from 11.45 to 7.95 on LLaMA-2 13B, demonstrating its effectiveness across model scales.
@@
For the coding domain-specific model, WizardCoder Luo et al. (2023), BitDistiller failed to offer the functional quantized models in both W3 and W2 settings. In our method, applying rotation alone was not effective in W2 settings and recovered some output quality in W3 settings. By incorporating LDP, we achieved up to a threefold improvement in performance, with accuracy increasing from 6.09% to 23.20% under the W2A4KV4. As shown in Fig. 9 with the application of LDP, we were able to produce logically correct code outputs and eliminate repetition of meaningless code generation.
For the mathematical reasoning model, MetaMath Yu et al. (2023), the baseline BitDistiller failed to offer functional quantized models while ours with LDP could produce working quantized models. These results highlight the critical role of LDP in enabling proper task performance for reasoning models under extreme low-bit quantization. The output comparison for this task is summarized in Fig. 10.
Long-Context Benchmarks
LLaMA-2 7B Method Avg.
Chat-4k BF16 32.53
BitDistiller 4.37
RCP 19.32
Instruct-32k BF16 27.13
BitDistiller 5.16
RCP 12.29
Table 4: Comparison of LongBench results of RCP under W2A4KV4 across different models and methods.
We conduct experiments on a subset of the LongBench dataset to evaluate the effectiveness of our method under various context lengths. Specifically, we test both LLaMA-2 7B-chat-4k and LLaMA-2 7B-Instruct-32k models across eight benchmark tasks. As shown in Table 4, our proposed RCP with W2A4KV4 consistently outperforms BitDistiller with W2A4KV4 across all tasks. For instance, on the LLaMA-2 7B-chat-4k model, RCP achieved an average score of 19.32, significantly higher than BitDistiller’s 4.37. Similarly, on the LLaMA-2 7B-Instruct-32k model, RCP yields 12.29 compared to BitDistiller’s 5.16, demonstrating robustness to extended context lengths. These findings further support the effectiveness of RCP-based quantization in preserving reasoning capability under constrained precision and longer context. The detailed results for each benchmark are presented in Table 15 and Table 16.
@@
Table 5 and 6 present the results for GEMV in terms of latency and memory comsumption. The latency of GEMV, excluding the activation quantization overhead, is faster compared to FP16 and QuaRot Ashkboos et al. (2024b). This improvement can be attributed to the lower bit precision, which enhances computational efficiency. Table 6 measures the peak memory footprint for W2A4 and W3A4. Although RCP incurs memory overhead due to additional parameters per quantization group beyond the BitDistiller, the performance gain from RCP compensates for this cost. For W2A4, a significant reduction on 5.29x in memory footprint was achieved compared to FP16. Note that in the LLaMA-3.2 series, it is necessary to separate the embedding table and head modules to satisfy the invariance arising from their tying. Furthermore, as the size of the embedding table has increased compared to previous models, the compression ratio has decreased accordingly.
The end-to-end inference latency is also measured under the W2-only setting, and the results are presented in Table 7. -BF16 is the original WizardCoder 7B, and -RCP is our model. We report prefill, decode, and total latency values. The weight-only version of our GEMV kernel is attached to the TinyChat engine Lin et al. (2024b) and several problems from the HumanEval benchmark where our RCP model provided correct answers are randomly chosen. We merged the non-uniform weight dequantization part of the W2A4 GEMV and the C++ PyTorch linear kernel (which works in BF16) into a single CUDA kernel to reduce launch overhead. The BF16 model used the C++ linear kernel only. In the prefill stage, RCP is slower than BF16 due to online dequantization overhead. In the decode stage, however, 2-bit weight representation with our efficient non-uniform dequantization greatly reduces the memory traffic from global to shared memory, resulting in an average decoding latency reduced by 30.19%. Combined, RCP can run faster on decoding-oriented tasks.
Problem #1 #2 #3 #4 #5 #6 #7 Avg.
Prefill-BF16 1.37 0.17 0.18 0.18 0.16 0.16 0.17 0.34
Prefill-RCP 1.29 0.23 0.24 0.25 0.23 0.19 0.22 0.38
Decode-BF16 13.18 13.35 13.45 13.08 13.25 13.83 13.52 13.38
Decode-RCP 9.26 9.32 9.47 9.15 9.08 9.81 9.27 9.34
Total-BF16 14.55 13.52 13.63 13.26 13.41 13.99 13.69 13.72
Total-RCP 10.55 9.55 9.71 9.40 9.31 10.00 9.49 9.72
Table 7: Comparison of end-to-end time per inference phase across different problems between BF16 and W2-only (LDP) quantization for WizardCoder 7B. All values are measured in ms/token.
Training Cost
Method Metric 1B 3B 7B 8B 13B
BitDistiller VRAM (GB) 35.1 42.2 32.0 77.4 130.2
Time (h) 64.0 64.8 68.4 93.6 29.6
Epoch 16 8 8 8 8
Batch 4 4 8 4 32
RCP VRAM (GB) 35.3 42.9 33.1 78.2 132.3
Time (h) 67.2 69.6 73.3 96.8 32.0
Epoch 16 8 8 8 8
Batch 4 4 8 4 32
Table 8: Comparison of VRAM and GPU usage for BitDistiller and RCP.
Table 8 summarizes the training configurations and training costs. The VRAM usage denotes the memory consumed on a single GPU and the GPU-hours were calculated by multiplying the training time by the total number of GPUs used. In our experiments, we conducted experiments on LLaMA-1, LLaMA-2 7B and LLaMA-3.2 (1B and 3B) on 8 RTX A6000 GPUs (48 GB each). For larger-scale models, LLaMA-3 8B was trained on 8 A100 GPUs (80GB each), and LLaMA-2 13B was trained on 8 GPUs (141GB each) within a DGX H200 system. The enlarged vocabulary in LLaMA-3 and later models increases gradient-computation demands, resulting in higher VRAM usage. To ensure training stability under these constraints, we set the training batch size to 4.
Our proposed RCP incurs approximately 10% additional training cost compared to the baseline. However, on LLaMA-2 7B, it improves perplexity from 17.40 to 8.31-nearly a twofold improvement. Furthermore, for LLaMA-3.2 3B, RCP improves the PPL by up to 42 times compared to the baseline. Considering these significant performance gains, the 10% additional training cost is acceptable.